# Heterogeneous Multi-Agent Reinforcement Learning for Zero-Shot Scalable Collaboration

### v1

@@ Abstract
The rise of multi-agent systems, especially the success of multi-agent reinforcement learning (MARL), is reshaping our future across diverse domains like autonomous vehicle networks. However, MARL still faces significant challenges, particularly in achieving zero-shot scalability, which allows trained MARL models to be directly applied to unseen tasks with varying numbers of agents. In addition, real-world multi-agent systems usually contain agents with different functions and strategies, while the existing scalable MARL methods only have limited heterogeneity. To address this, we propose a novel MARL framework named Scalable and Heterogeneous Proximal Policy Optimization (SHPPO), integrating heterogeneity into parameter-shared PPO-based MARL networks. we first leverage a latent network to adaptively learn strategy patterns for each agent. Second, we introduce a heterogeneous layer for decision-making, whose parameters are specifically generated by the learned latent variables. Our approach is scalable as all the parameters are shared except for the heterogeneous layer, and gains both inter-individual and temporal heterogeneity at the same time. We implement our approach based on the state-of-the-art backbone PPO-based algorithm as SHPPO, while our approach is agnostic to the backbone and can be seamlessly plugged into any parameter-shared MARL method. SHPPO exhibits superior performance over the baselines such as MAPPO and HAPPO in classic MARL environments like Starcraft Multi-Agent Challenge (SMAC) and Google Research Football (GRF), showcasing enhanced zero-shot scalability and offering insights into the learned latent representation’s impact on team performance by visualization.
@@
Multi-agent reinforcement learning (MARL), artificial intelligence, multi-agent system, heterogeneity, scalability, representation learning.
@@ I Introduction
Multi-agent systems are playing more and more important roles in future human society with the development of artificial intelligence [1, 2, 3, 4, 5]. Multi-agent reinforcement learning (MARL) has seen widespread application in cooperative multi-agent tasks across various domains like autonomous vehicle networks [6, 7, 8], guaranteed display ads [9], and video games [10], significantly enhancing collaboration among agents. With the introduction of actor-critic [11] to MARL, recent advances [12, 13] in MARL algorithms exhibit great potential for addressing more complex cooperative tasks in dynamic environments. Despite its successes, traditional MARL methods are confined to specific environments where models are trained. Lack of generalizability poses a significant barrier to real-world deployment.
Specifically, in this paper, we focus on zero-shot scalability and the generalizability of agents’ population sizes. This involves directly applying trained MARL models to other unseen tasks with similar settings but varying numbers of agents without the need for further training, which is necessary in various real-world applications. For instance, traffic flows fluctuate throughout the day, requiring a traffic junction coordination algorithm capable of transferring to scenarios with varying numbers of vehicles.
@@
Figure 1: Illustration of scalability and heterogeneity in SMAC. (a) The original task without heterogeneity, where all the agents have the same strategy. (b) The original task with heterogeneity, where the agents form two groups to attract the fire (blue) and attack in the distance (red). (c) The unseen task with a new agent without heterogeneity, where the new agent has the same strategy. (d) The new task with heterogeneity, where the new agent can adaptively choose the strategy.
@@
Moreover, agents within a team may have different strategies or even different functions; thus, learning and transferring heterogeneous roles in a team stands out as an important problem of zero-shot scalable collaboration. First, lack of heterogeneity hinders agents from effectively collaborating with diverse counterparts, constraining team performance. As shown in Fig. 1(a), when heterogeneity is not considered, the agents will have similar strategies and simply form a circle to attack in the game screenshot, while there are different kinds of units with different capacities in the team. But in Fig. 1(b), after the introduction of heterogeneity, some agents may attack in the front to attract the fire (blue circle) while others stay back to be covered. Moreover, in zero-shot scalable tasks, teamwork roles change with team sizes (refer to our experiments in Fig. 8). An adaptive policy is needed to flexibly assign and adjust agents’ heterogeneous roles for better zero-shot scalability. We can implement this policy based on heterogeneous settings (Fig. 1(d)) but may fail without heterogeneity (Fig. 1(c)).
@@
The literature has attempted to address heterogeneity and scalability issues independently. Some works [14, 15] explore giving each agent a distinct model to improve heterogeneity, but they struggle to scale with the number of agents. They suffer from the dimension curse, and there are no training processes for the newly added agents. While other works [16, 17] try to build population-invariant MARL for scalability but ignore the heterogeneity. In addition, parameter sharing is scalable and widely used as network parameters for all agents are shared, bringing benefits of training efficiency and superior performance [18, 19, 20], which, however, also limit the diversity and temporal flexibility of the policies. Motivated by these challenges, we aim to enhance zero-shot scalable collaboration by integrating both inter-individual and temporal heterogeneity.
@@
In this paper, we propose a novel MARL framework called Scalable and Heterogeneous Proximal Policy Optimization (SHPPO) to add a heterogeneous layer to a parameter-shared version of the classic PPO [21] under multi-agent settings (see Fig. 2). We introduce a latent network to adaptively learn low-dimension latent variables to represent the strategy pattern for each agent, according to its observations and trajectories. And then, based on the latent variable, SHPPO generates the parameters of the heterogeneous layer in the actor network. We also introduce a centralized inference net to guide the learning of the latent network to form a symmetrical structure as the actor-critic in PPO (see Fig. 2(b)). Note that our approach can be applied to any parameter-shared MARL backbone, and we take PPO as an example to evaluate the performance in this paper.
@@
In this way, though all the network parameters except for the heterogeneous layers are shared for scalability, the heterogeneous layer for every agent is specifically designed to enhance the diversity of the team. In addition to the inter-individual heterogeneity, the agents are able to update the strategy with the change of observations during the task, so the agents also gain temporal heterogeneity. When transferred to a new scenario with new agents added or removed, the latent variable is accordingly updated with the learned labor division knowledge. Thus the following generation of new heterogeneous layers can transfer the learned strategy patterns while keeping heterogeneity.
@@ II-A MARL for Collaboration
Multi-Agent Reinforcement Learning (MARL) has emerged as a powerful paradigm for training groups of AI agents to collaborate and solve complex tasks [6, 7, 8, 9, 10]. QMIX [24] addresses the challenge of credit assignment by computing joint Q-values from individual agents as a value-based method, while MADDPG [12] extends the policy gradient method DDPG [25] to multi-agent settings using centralized critics and decentralized actors. Different from the off-policy MADDPG, MAPPO [20], based on on-policy PPO [21], achieves strong performance in various MARL benchmarks. Despite significant progress, challenges in MARL persist, including generalizability for unseen scenarios, sample efficiency, non-stationarity, and agent communication [26, 27, 28, 29]. In this paper, we mainly focus on zero-shot scalability as one of the generalizability issues.
@@ III-A DEC-POMDP
In the development of multi-agent reinforcement learning, various paradigms have been introduced to adapt single-agent training methodologies. Initially, the Decentralized Training with Decentralized Execution (DTDE) [46] paradigm is employed, where each agent autonomously makes decisions and learns independently, operating without access to the observations, actions, or policies of other agents. However, this approach encounters the non-stationary problem from the perspective of every single agent when all agents individually update their policies.
To enhance the stability of the learning process, the Centralized Training with Centralized Execution (CTCE) paradigm was introduced [47]. This method conceptualizes the joint policy as a centralized policy and utilizes single-agent methods to learn this aggregated policy. Nonetheless, scalability is hampered by the curse of dimensionality, as the state-action space grows exponentially with the number of agents.
To address this scalability issue, the Centralized Training with Decentralized Execution (CTDE) paradigm [12] combines the advantages of both DTDE and CTCE. CTDE allows agents to utilize additional information during training, yet during execution, each agent acts independently, and decision-making relies solely on its own policy.
When agents are unable to observe the concise state of the environment, commonly referred to as partial observation (PO), a Partially Observable Markov Decision Process (POMDP) extends the MDP model by incorporating observations and their conditional probability of occurrence based on the state of the environment [48, 49]. Within the POMDP and CTDE paradigm, the decentralized partially observable Markov decision processes (DEC-POMDP) [50] have become a widely adopted model in MARL. A DEC-POMDP is defined by $<N,S,A,O,r,P,\gamma>$. $o_{i}=O(s;i)$ represents the local observation for agent $i$ at the global state $s$. Each agent utilizes its policy $\pi_{\theta_{i}}(a_{i}|o_{i})$, parameterized by $\theta_{i}$, to produce an action $a_{i}$ from the local observation $o_{i}$. And agents jointly aim to optimize the discounted accumulated reward $J(\theta)=\mathbb{E}_{\boldsymbol{a},s}[\sum_{t}\gamma^{t}r(s,\boldsymbol{a})]$, where $\boldsymbol{a}$ is the joint action at time step $t$.
Figure 2: Schematics of our approach SHPPO. (a) Architecture of LatentNet. (b) Framework of SHPPO. (c) Architecture of ActorNet with the heterogeneous layer (HeteLayer). (d) Architecture of HeteLayer. The pink blocks denote the losses. The yellow and green bars denote different HeteLayers.
@@ IV Scalable and Heterogeneous
Existing heterogeneous MARL methods such as HAPPO model each agent as an independent and different network; it suffers from its inflexibility and huge parameter costs when the scale of the multi-agent system increases. In this section, we propose a novel MARL framework SHPPO and its training method to tackle the trade-off of scalability and heterogeneity.
We try to solve this problem in another direction different from HAPPO: we add adaptive heterogeneity to one single parameter-shared model. Specifically, as in Fig. 2(b), we design the new framework as two parts, latent learning and action learning. First, we implement a latent learning model (LatentNet) to adaptively generate a latent distribution $\mathcal{D}$ for each agent based on its observation and memory (in practice, we take the hidden state of the RNN in the actor network (ActorNet) as the memory). The learned latent distribution represents the strategy pattern of the specific agent, which is further sampled and applied to the computation of the heterogeneous layer parameters for this agent. Except for the heterogeneous layer, the other layers’ parameters in the ActorNet are shared by all the agents. When the number of agents varies, the LatentNet can always generate a unique latent distribution for each agent so that the proposed method can handle the scalability and heterogeneity at the same time. Our approach is agnostic to the implementation of the action learning part, and in practice, we adopt the shared version of HAPPO as the backbone for its monotonic improvement property.
@@ IV-A Scalable Latent Learning
To design a heterogeneous layer for each agent, we first need to represent the agent’s strategy pattern as a low-dimensional latent variable $l$ as shown in Fig. 2(a). Take agent $i$ as an example, we take a 3-layer MLP as the Encoder to convert the observation $o_{i}$ and the memory from the RNN $h^{t-1}_{i}$ at time step $t$ to a latent distribution. Therefore, the Encoder outputs the mean $\boldsymbol{\mu}_{i}$ and the standard deviation $\boldsymbol{\sigma}_{i}$ of the multi-dimensional Gaussian $\boldsymbol{\mathcal{D}}_{i}$. Like humans, the agents’ strategy patterns may not be exactly the same under similar situations, so the Encoder learns a distribution instead of a vector. The final latent variable $l_{i}$ is generated by sampling to keep the randomness as
@@
Inspired by the classic Actor-Critic architecture [11], an inference net (InferenceNet) is introduced as a critic to guide the learning of the LatentNet (see Fig. 2(b)). Though the LatentNet generates the latent variable individually for each agent, the InferenceNet centrally learns an overall value function to evaluate the whole team. Specifically, the InferenceNet takes all the $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$ as the inputs, together with the joint global observation $o_{g}$, and then computes $V_{I}$ as the evaluation for the LatentNet. The InferenceNet is trained by supervised learning to minimize the difference between $V_{I}$ and the return from the environment $R$. In another way, it judges whether the learned latent distributions are good enough to help with the decision-making based on global observations. Thus, the LatentNet’s goal is to maximize the $V_{I}$.
It is noteworthy that all the agents share the same parameters. Thanks to the CTDE architecture, during the execution and transfer, we only forward the LatentNet and ActorNet, namely, the Agent part in gray color in Fig. 2(b). Though the input dimensions of the InferenceNet and CriticNet vary with the population sizes, this will not influence the transfer to the new scenario to execute. In this way, our method is scalable to fit scenarios with different agent numbers.

### v2

@@ Abstract
The emergence of multi-agent reinforcement learning (MARL) is significantly transforming various fields like autonomous vehicle networks. However, real-world multi-agent systems typically contain multiple roles, and the scale of these systems dynamically fluctuates. Consequently, in order to achieve zero-shot scalable collaboration, it is essential that strategies for different roles can be updated flexibly according to the scales, which is still a challenge for current MARL frameworks. To address this, we propose a novel MARL framework named Scalable and Heterogeneous Proximal Policy Optimization (SHPPO), integrating heterogeneity into parameter-shared PPO-based MARL networks. We first leverage a latent network to learn strategy patterns for each agent adaptively. Second, we introduce a heterogeneous layer to be inserted into decision-making networks, whose parameters are specifically generated by the learned latent variables. Our approach is scalable as all the parameters are shared except for the heterogeneous layer, and gains both inter-individual and temporal heterogeneity, allowing SHPPO to adapt effectively to varying scales. SHPPO exhibits superior performance in classic MARL environments like Starcraft Multi-Agent Challenge (SMAC) and Google Research Football (GRF), showcasing enhanced zero-shot scalability, and offering insights into the learned latent variables’ impact on team performance by visualization.
Note to Practitioners
Multi-agent reinforcement learning (MARL) algorithms are extensively utilized in cooperative multi-agent scenarios, such as multiple rescue unmanned aerial vehicles (UAVs), autonomous vehicles navigating crowded intersections efficiently, and robots jointly managing cargo at ports. However, real-world applications necessitate both scalable and heterogeneous collaboration strategies. For instance, rescue UAV teams may comprise varying numbers of units designated for searching, firefighting, carrying, and communication, depending on the specific tasks. To address this, our approach, SHPPO, can adaptively reorganize roles within the team without requiring additional training when team sizes fluctuate. This superior zero-shot scalability of heterogeneous strategies enables the learned MARL models to be effectively applied to more complex and dynamic tasks. Although a notable limitation of MARL is the disparity between current virtual training environments and real-world scenarios, we are optimistic that advancements in simulation and modeling technologies will soon bridge this gap.
@@
Multi-agent reinforcement learning, artificial intelligence, multi-agent system, heterogeneity, scalability.
@@ I Introduction
Figure 1: Illustration of scalability and heterogeneity in SMAC. (a) The original task without heterogeneous strategies, where different kinds of agents have the same strategy as a circle. (b) The original task with heterogeneous strategies, where the agents form two groups to attract the fire (blue) and attack in the distance (red). (c) The unseen task with a new agent (green) without heterogeneous strategies, where the new agent has the same strategy as others. (d) The new task with heterogeneous strategies, where the new agent can adaptively choose the strategy.
Multi-agent systems are playing more and more important roles in future human society with the development of artificial intelligence [1, 2, 3, 4, 5]. Multi-agent reinforcement learning (MARL) has seen widespread application in cooperative multi-agent tasks across various domains like autonomous vehicle networks [6, 7, 8], guaranteed display ads [9], and video games [10], significantly enhancing collaboration among agents. With the introduction of actor-critic [11] to MARL, recent advances [12, 13] in MARL exhibit great potential for addressing more complex cooperative tasks.
Despite its successes, traditional MARL methods are confined to specific environments where models are trained. However, in real-world applications, the tasks are usually dynamic and diverse, with issues of scalability (variation of the number of agents) and heterogeneity (diversity of agents’ functions and abilities), while it is usually hard to balance the trade-off of scalability and heterogeneity. Simpler mechanisms tend to generalize more effectively across different scales; however, this often comes at the expense of the specificity required for heterogeneous strategies. In such cases, vanilla MARL methods may fail to transfer the learned strategy for a heterogeneous multi-agent system to a new scenario with agents added or removed, as the role assignment and the collaboration mechanism will change. Thus, both scalability and heterogeneity should be considered to design a more practical MARL algorithm.
Specifically, in this paper, we focus on zero-shot scalability. This involves directly applying trained MARL models to other unseen tasks with similar settings but varying numbers of agents without the need for further training, which is necessary for various real-world applications. For instance, traffic flows fluctuate throughout the day, requiring a traffic junction coordination algorithm capable of transferring to scenarios with varying numbers of vehicles.
@@
Moreover, learning and transferring heterogeneous roles in a team stands out as an important problem of zero-shot scalable collaboration. First, lack of heterogeneity hinders agents from effectively collaborating with diverse counterparts, constraining team performance. As shown in Fig. 1(a), when heterogeneity is not considered, the agents will have similar strategies and simply form a circle to attack in the game screenshot, while there are different kinds of units with different capacities in the team. But in Fig. 1(b), after the introduction of heterogeneity, some agents may attack in the front to attract the fire (blue circle) while others stay back to be covered. Second, in unseen tasks with a varied number of agents, the role assignment changes with team sizes (refer to our experiments in Fig. 6). An adaptive policy is needed to flexibly assign and adjust agents’ heterogeneous roles for better zero-shot scalability. We can implement such a policy based on heterogeneous strategies (Fig. 1(d)), but may fail without heterogeneity (Fig. 1(c)). Therefore, it is significant to answer the question: "How to design a novel MARL framework learning adaptive heterogeneous strategies to achieve better zero-shot scalability?", while the literature only attempted to address the two issues independently.
@@
Some works [14, 15] explore giving each agent a distinct model to improve heterogeneity, but they struggle to scale with the number of agents. They suffer from the dimension curse, and there are no training processes for the newly added agents. While other works [16, 17] try to build population-invariant MARL for scalability but ignore the heterogeneity. In addition, parameter sharing is scalable and widely used as network parameters for all agents are shared, bringing benefits of training efficiency and superior performance [18, 19, 20], which, however, also limit the diversity and temporal flexibility of the policies. Motivated by these challenges, we aim to enhance zero-shot scalable collaboration by integrating both inter-individual (the agents’ strategies differ from each other) and temporal (the agents’ strategies update with the progress) heterogeneity.
@@
In this paper, we propose a novel MARL framework called Scalable and Heterogeneous Proximal Policy Optimization (SHPPO) to add a heterogeneous layer to a parameter-shared version of PPO [21] under multi-agent settings (see Fig. 2). We introduce a latent network to adaptively learn low-dimension latent variables to represent the strategy pattern for each agent, according to its observations and trajectories. Then, based on the latent variable, SHPPO generates the parameters of the heterogeneous layer in the actor network. We also introduce a centralized inference net to guide the learning of the latent network to form a symmetrical structure as the actor-critic in PPO (see Fig. 2(b)). Note that our approach can be applied to any parameter-shared MARL backbone, and we take PPO as an example to evaluate the performance in this paper.
@@
In this way, though all the network parameters except for the heterogeneous layers are shared for scalability, the heterogeneous layer for every agent is specifically designed to enhance the diversity of the team. In addition to the inter-individual heterogeneity, the agents can update the strategy with the change of observations during the task, so the agents also gain temporal heterogeneity. When transferred to a new scenario with new agents added or removed, the latent variables are accordingly updated with the learned labor division knowledge (See Fig. 6). Thus the following generation of new heterogeneous layers can transfer the learned strategy patterns to achieve zero-shot scalability while keeping heterogeneity.
@@ II-A MARL for Collaboration
Multi-Agent Reinforcement Learning (MARL) has emerged as a powerful paradigm for training groups of AI agents to collaborate and solve complex tasks [6, 7, 8, 9, 10, 24, 25, 26]. QMIX [27] addresses the challenge of credit assignment by computing joint Q-values from individual agents as a value-based method, while MADDPG [12] extends the policy gradient method DDPG [28] to multi-agent settings using centralized critics and decentralized actors. Different from the off-policy MADDPG, MAPPO [20], based on on-policy PPO [21], achieves strong performance in various MARL benchmarks. Despite significant progress, challenges in MARL persist, including generalizability for unseen scenarios, sample efficiency, non-stationarity, and agent communication [29, 30, 31, 32]. In this paper, we mainly focus on zero-shot scalability as one of the generalizability issues.
@@ III-A DEC-POMDP
However, in naive MARL, scalability is hampered by the curse of dimensionality, as the state-action space grows exponentially with the number of agents. Besides, there is the non-stationary problem when agents individually update their policies. To address these two issues, the Centralized Training with Decentralized Execution (CTDE) paradigm [12] is proposed. CTDE allows agents to utilize central information during training, yet during execution each agent acts independently.
When agents are unable to observe the concise state of the environment, commonly referred to as partial observation (PO), a Partially Observable Markov Decision Process (POMDP) extends the MDP model by incorporating observations and their conditional probability of occurrence based on the state of the environment [49, 50]. Within the POMDP and CTDE paradigm, the decentralized partially observable Markov decision processes (DEC-POMDP) [51] have become a widely adopted model in MARL. A DEC-POMDP is defined by $<N,S,A,O,r,P,\gamma>$. $o_{i}=O(s;i)$ represents the local observation for agent $i$ at the global state $s$. Each agent utilizes its policy $\pi_{\theta_{i}}(a_{i}|o_{i})$, parameterized by $\theta_{i}$, to produce an action $a_{i}$ from the local observation $o_{i}$. And agents jointly aim to optimize the discounted accumulated reward $J(\theta)=\mathbb{E}_{\boldsymbol{a},s}[\sum_{t}\gamma^{t}r(s,\boldsymbol{a})]$, where $\boldsymbol{a}$ is the joint action.
@@ IV Scalable and Heterogeneous
Figure 2: Schematics of our approach SHPPO. (a) Architecture of LatentNet to represent the agent’s strategy pattern as a low-dimensional latent variable $l$. (b) Framework of SHPPO to conduct latent learning and action learning with dual actor-critic networks. Each agent contains two networks ActorNet and LatentNet. The two networks are parameter-shared among the agents except for the heterogeneous layer (HeteLayer), so SHPPO is scalable. (c) Architecture of ActorNet with HeteLayer, bringing heterogeneity in SHPPO. (d) Architecture of HeteLayer where the latent variable $l$ is decoded to the parameters of the heterogeneous linear layer. The pink blocks denote the losses. The yellow and green bars denote different HeteLayers.
Existing heterogeneous MARL methods, such as HAPPO, model each agent as an independent and different network; it suffers from its inflexibility and huge parameter costs when the scale of the multi-agent system increases. Especially, when new agents are introduced, there is no pre-trained model available for them, as each model is unique. In this section, we propose a novel MARL framework SHPPO, which can learn adaptive heterogeneous strategies that can be transferred for zero-shot scalable collaboration.
We try to solve this problem in another direction different from HAPPO: we add adaptive heterogeneity to one single parameter-shared model. Specifically, as in Fig. 2(b), we structure the new framework into two components: latent learning and action learning. First, we implement a latent learning model (LatentNet) to adaptively generate a latent distribution $\mathcal{D}$ for each agent based on its observation and memory (in practice, we take the hidden state of the RNN in the actor network (ActorNet) as the memory). The learned latent distribution represents the strategy pattern of the specific agent, which is further sampled and applied to the computation of the heterogeneous layer parameters for this agent. Except for the heterogeneous layer, the other layers’ parameters in the ActorNet are shared by all the agents. When the number of agents varies, the LatentNet can always generate a unique latent distribution for each agent so that the proposed method can handle the scalability and heterogeneity at the same time. Our approach is agnostic to the implementation of the action learning part, and in practice, we adopt the shared version of HAPPO as the backbone for its monotonic improvement property.
In the following subsections, we first introduce the latent learning (Section IV-A) and heterogeneous layers (Section IV-B). Then, we elaborate on the loss design (Section IV-C) and the efficacy of the mechanism (Section IV-D). Finally, we conclude with the overall training procedure (Section IV-E).
@@ IV-A Scalable Latent Learning
To design a heterogeneous layer for each agent, we first need to represent the agent’s strategy pattern as a low-dimensional latent variable $l$ as shown in Fig. 2(a). Take agent $i$ as an example, we take a 3-layer MLP as the Encoder to convert the observation $o_{i}$ and the memory from the RNN $h^{t-1}_{i}$ at time step $t$ to a latent distribution. Specifically, the Encoder outputs the mean $\boldsymbol{\mu}_{i}$ and the standard deviation $\boldsymbol{\sigma}_{i}$ of the multi-dimensional Gaussian $\boldsymbol{\mathcal{D}}_{i}$. Like humans, the agents’ strategy patterns may not be exactly the same under similar situations, so the Encoder learns a distribution instead of a fixed vector. The final latent variable $l_{i}$ is generated by sampling to keep the randomness as
@@
Inspired by the classic Actor-Critic architecture [11], an inference net (InferenceNet) is introduced as a critic to guide the learning of the LatentNet (see Fig. 2(b)). Though the LatentNet generates the latent variable individually for each agent, the InferenceNet centrally learns an overall value function to evaluate the whole team. Specifically, the InferenceNet takes $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$ from all the agents as the inputs, together with the joint global observation $o_{g}$, and then computes $V_{I}$ as the evaluation for the LatentNet. The InferenceNet is trained by supervised learning to minimize the difference between $V_{I}$ and the return from the environment $R$. In another way, it judges whether the learned latent distributions are good enough to help with decision-making based on global observations. Thus, the LatentNet’s goal is to maximize the $V_{I}$.
It is noteworthy that all the agents share the same parameters. Thanks to the CTDE architecture, during the execution and transfer, we only forward the LatentNet and ActorNet, namely, the Agent part in gray color in Fig. 2(b). Though the input dimensions of the InferenceNet and CriticNet vary with the population sizes, this will not influence the transfer to the new scenario to execute. In this way, our method is scalable to fit scenarios with different agent numbers. Please refer to Appendix Table IV for the detailed network settings.